# Hiba Arnaout1, Noy Sternlicht2, Tom Hope2,3, Iryna Gurevych1

### v1

@@ F.1 A new dataset for identifying impact-revealing citations
Since our task on classifying citation context into “impact-revealing” (for both confirmation/praise and correction/critique) or “other” is new, there is a lack of groundtruth data. To construct such a resource, we build upon on an existing human annotated dataset, namely PST-Bench Zhang et al. (2024). An instance in this dataset consists of a pair of research papers annotated for whether one of them influenced (impacted) the other. Influence here is defined only through a positive lens, and is further restricted to the following two intents: “inspiration” and “motivation”. To create our dataset, first, we randomly select 1k influential (impact-revealing) papers, as labeled by the PST-Bench dataset annotators, and look up their aggregated citation contexts [^12] The cited paper might be mentioned several times in the citing paper.. Since the meaning of influential, according to PST-Bench, only focuses on two intents “inspiration” and “motivation”, both limited to praise, we augment the dataset with other types for both confirmatory and correction citation intents. To do so, we manually craft a handful of textual patterns such as “has led to open questions/challenges/unresolved issues” to capture intents indicating research limitations, then ask GPT4-o to provide a longer list with variations of these patterns. We end up with a total of 33 patterns (examples shown in Figure 12). Following this step, we crawl $\sim$ 200k aggregated citation contexts and search them using the list of impact-revealing phrases, then randomly sample 1k instances from the matching cases. To augment the data with non-impact-revealing (“other”) examples, we sample 2k non-influential instances from the PST-Bench dataset. These are citations that are listed under the category “other citations”, as in non-influential ones. To ensure the deemed non-influential examples are not in fact impact-revealing (ones which belong to intents other than “motivation” and “inspiration”), we verify that there is no match with all the impact-revealing phrases. Our newly constructed dataset, which we release with this work, consists of 4k citation contexts augmented with either “impact-revealing” or “other” intent class (balanced).
@@ F.2 Details about comparisons with intent classifiers
Sample results are shown in Table F.2.
@@ G.1 Examples
G.2 Results per field
@@ G.4 G-Eval reasoning
G.5 Visualized impact statements
@@
Examples are shown in Figures 13 (computer science paper),  14 (psychology paper),  15 (psychology paper), and  16 (medicine paper). These impact summaries align closely with our definition. They capture the evolution of citations intents, from initial adoption and theoretical contributions to critiques, methodological refinements, policy influences, and integration into modern applications. For example, the impact summary of “Hierarchically Classifying Documents Using Very Few Words” (ICML’97) illustrates how the paper initially shaped the research paradigm in document classification across fields like bio-informatics and legal document analysis. Later, it faced critiques related to scalability and multi-class accuracy, leading to refining the existing method. Another example is the impact summary of “Sex Bias in Neuroscience and Biomedical Research” (Neuroscience & Biobehavioral Reviews’11), which highlights the early discussions on gender disparities that this paper sparked. Those discussions later led to policy changes, promoting balanced gender representation in preclinical studies, influencing new research methods across various disciplines.
Appendix H More on the human evaluation
H.1 Participants
Nationalities: 2 German, 2 British, 2 Chinese, 1 American, 1 Czech, 1 Albanian; genders: 5 male, 4 female; research areas: AI, NLP, knowledge graphs, databases and information systems, computational social sciences, psychology and brain sciences.
H.2 Perceived usefulness
Results are in Figure 17. We show the overall results in (a), and results for two subsets of papers, namely papers with the most impact-revealing citations in (b), and papers with the highest number of citations in (c). Both subsets show impact summaries that are rated higher, likely because higher counts of impact-revealing citations and citations in general allow for richer contexts, which provide the model with clearer signals about the paper’s impact, leading to more accurate and insightful summaries.
H.3 Open-feedback
A couple professors found certain summaries too generic, failing to highlight specific strengths and limitations. They also pointed out that summaries based on citations from lesser-known conferences may not fully reflect a paper’s true influence, and the quality of summaries often correlates with the size of impact-revealing citation context (a point further validated in Figure 17 (b)). Additionally, one professor noted that for some papers, it was difficult to assess impact because the work was collaborative, and they were only familiar with certain aspects of the paper’s contributions, while their co-authors might be more familiar with other aspects. Other professors suggested improvements in both content and structure: they recommended incorporating citation counts within the text, which we plan to implement in the future, and enhancing the structure of our semi-structured summaries by adding elements like bullet points to improve readability.
@@ I.1 Equal citation count, different citing reasons
We select 10 diverse research topics (e.g., LLMs for code generation, cultural sensitivity in clinical psychology), and analyze 3 papers within each. We ensure that the topic-specific papers have close citation counts. In this use case, we demonstrate how fine-grained citation intent analysis provides a deeper understanding of research impact, moving beyond raw citation counts. Results are in Tables 17 and 18. We observe that some papers’ main impact is inspiring or motivating new work, some are used for their method or data, others are cited to point out remaining challenges and propose improvements, etc. For example, for papers on open-information extraction with citation counts of $\sim$ 200,  Cui et al. (2018) is frequently cited to motivate new work or highlight existing limitations, while  Gashteovski et al. (2017) and  Han et al. (2019) are cited for their method. In commonsense knowledge mining, only one of the 3 papers is often cited for its dataset Hwang et al. (2021), and even though the other 2 papers are cited to highlight limitations, the limitations are different, namely challenges with handling constraints and reasoning for one Davison et al. (2019) , but data noise and limited coverage for the other Speer and Havasi (2013). These examples show that even when citation counts are similar, citation intents reveal the specific citing reasons, shaping the perception of a paper’s true impact and assisting in the impact summary generation task.
@@ I.2 Author-level impact summaries
F1 agreement 0.83 0.71 0.66 0.86 0.74
Table 12: Human and LLM correlation and agreement on evaluating results (Section 3.3). p-value: $\leq$0.001, *$\leq$0.05.
@@
Table 14: Examples of impact summaries from the informativeness evaluation, baseline = zero-shot variant, our method = impact-revealing variants.
@@
Figure 17: Perceived usefulness results for: (a) all papers, (b) papers in the top 10% for number of impact-revealing citations, (c) papers in the top 10% for total number of citations.

### v2

@@ F.1 A new dataset for identifying impact-revealing citations
Since our task on classifying citation context into “impact-revealing” (for both confirmation/praise and correction/critique) or “other” is new, there is a lack of groundtruth data. To construct such a resource, we build upon on an existing human annotated dataset, namely PST-Bench Zhang et al. (2024a). An instance in this dataset consists of a pair of research papers annotated for whether one of them influenced (impacted) the other. Influence here is defined only through a positive lens, and is further restricted to the following two intents: “inspiration” and “motivation”. To create our dataset, first, we randomly select 1k influential (impact-revealing) papers, as labeled by the PST-Bench dataset annotators, and look up their aggregated citation contexts [^13] The cited paper might be mentioned several times in the citing paper.. Since the meaning of influential, according to PST-Bench, only focuses on two intents “inspiration” and “motivation”, both limited to praise, we augment the dataset with other types for both confirmatory and correction citation intents. To do so, we manually craft a handful of textual patterns such as “has led to open questions/challenges/unresolved issues” to capture intents indicating research limitations, then ask GPT4-o to provide a longer list with variations of these patterns. We end up with a total of 33 patterns (examples shown in Figure 13). Following this step, we crawl $\sim$ 200k aggregated citation contexts and search them using the list of impact-revealing phrases, then randomly sample 1k instances from the matching cases. To augment the data with non-impact-revealing (“other”) examples, we sample 2k non-influential instances from the PST-Bench dataset. These are citations that are listed under the category “other citations”, as in non-influential ones. To ensure the deemed non-influential examples are not in fact impact-revealing (ones which belong to intents other than “motivation” and “inspiration”), we verify that there is no match with all the impact-revealing phrases. Our newly constructed dataset, which we release with this work, consists of 4k citation contexts augmented with either “impact-revealing” or “other” intent class (balanced). After manually inspecting a random sample of 100 instances, we found that 90% were correctly identified as impact-revealing; the remaining 10% included rare cases where phrases like motivated by” appeared to signal impact (e.g., Voters are motivated by partisan social identities… (Greene 2004)”) but did not reflect the actual citation intent.
@@ F.2 Details about comparisons with intent classifiers
Sample results are shown in Table 11.
@@ G.1 Examples
G.2 Long tail impact coverage
We conduct a focused, quantitative analysis to evaluate the model’s ability to capture long-tail yet meaningful aspects of a paper’s impact. Specifically, we manually select 10 intent themes associated with 10 different papers that fall outside the top-3 most frequent themes (and are therefore not covered in the Coverage@3 metric) but nonetheless reflect important impact dimensions. We then analyze 50 generated summaries, 10 per variant, and measure the extent to which these summaries capture the selected long-tail themes. We refer to this metric as long tail impact coverage, or LTI for short. As shown in Table 12, LTI ranges from 30% for the “all citation context, no intents” variant to 50% when citation intents are included. These findings suggest that the model is capable of recognizing and reflecting impact-revealing citations even when it appears less frequently in the input text, highlighting its sensitivity to semantically important but infrequent signals.
Prompt variant LTI Coverage
Citations Intents
None 0.1
All 0.3
All 0.5
Impact-rev. 0.4
Impact-rev. 0.4
Table 12: LTI: long tail impact coverage
G.3 Results per field
@@ G.5 G-Eval reasoning
G.6 Visualized impact statements
@@
Examples are shown in Figures 14 (computer science paper),  15 (psychology paper),  16 (psychology paper), and  17 (medicine paper). These impact summaries align closely with our definition. They capture the evolution of citations intents, from initial adoption and theoretical contributions to critiques, methodological refinements, policy influences, and integration into modern applications. For example, the impact summary of “Hierarchically Classifying Documents Using Very Few Words” (ICML’97) illustrates how the paper initially shaped the research paradigm in document classification across fields like bio-informatics and legal document analysis. Later, it faced critiques related to scalability and multi-class accuracy, leading to refining the existing method. Another example is the impact summary of “Sex Bias in Neuroscience and Biomedical Research” (Neuroscience & Biobehavioral Reviews’11), which highlights the early discussions on gender disparities that this paper sparked. Those discussions later led to policy changes, promoting balanced gender representation in preclinical studies, influencing new research methods across various disciplines.
G.7 Error analysis of our best variant
Although our intent classifier outperforms established baselines (Section 3.2), misclassifications at the intent prediction stage can still compromise the quality of the generated impact summaries. For example, when an incidental citation context is mistakenly classified as impact-revealing, the resulting summary may include statements that do not reflect the true scholarly impact. For instance, in the following part of an impact summary, “impact period: 2003-2010 Initial Contextual Reference: During this early period post-publication, the paper was often cited to provide comprehensive background information on the prevalence.. and risks associated with Alzheimer’s and Parkinson’s diseases…”, we can see that being mentioned for background information and context is listed under one of the impact periods. This is not in line with our definition of impact summaries, which only restrict impact to direct use of work. We inspect the reason for this mistake and track it back to a misclassification in the previous step, where intents such as “providing context on the prevalence and significance of Parkinson’s disease” were classified as impact-revealing.
Conversely, when an impact-revealing citation is misclassified as incidental, the summary is deprived of key evidence of impact. For instance, we observed that a statement in the impact summary describing the use of a graph neural network methodology for drug discovery disappears when the related citation contexts, those referencing this methodological application, are removed from the input.
These cases illustrate the sensitivity of the generation step to errors in intent classification, underscoring the importance of continued research in this area. We hope our work lays the groundwork for future studies aimed at improving intent classification as a critical component of generating accurate and informative impact summaries.
G.8 Generating impact summaries using other LLMs
We selected GPT-4o as our primary language model for this task due to its state-of-the-art performance in long-context reasoning and summarization. While we recognize the potential biases that may arise from relying solely on a single model family, the use of a consistent LLM and evaluation framework enables us to more clearly isolate the effects of input design. To evaluate the robustness of our method across model families, we also generate impact summaries using Qwen-2.5-72B Yang et al. (2025) and Gemini-2.5-flash Comanici et al. (2025), in addition to GPT-4o.
Quantitative results are presented in Table 13. Interestingly, the baseline variant using Qwen achieves the highest insightfulness score. However, this comes at the cost of faithfulness, with the baseline producing 18% fewer faithful claims than our best-performing variant. In contrast, Gemini demonstrates strong performance in citation year compliance, particularly with the impact-revealing with intents variant. Gemini tends to include numerous citations to substantiate its claims during impact periods and accurately aligns each citation with its corresponding time period. This variant also achieves the highest faithfulness (0.96, tied with all citations with intent variant), coverage (0.58), and specificity (0.83).
We present example summaries generated by both Qwen and Gemini in Tables 14 and  15.
Prompt variant Trustworthiness Informativeness
Citations Intents Faith. Cov. Cyc. Insi. Trend. Spec.
Qwen
None 0.70 0.42 n/a 0.73 0.84 0.76
All 0.86 0.44 0.46 0.67 0.81 0.78
All 0.86 0.44 0.44 0.66 0.80 0.78
Impact-rev. 0.86 0.44 0.43 0.70 0.81 0.77
Impact-rev. 0.88 0.45 0.42 0.68 0.81 0.76
Gemini
None 0.91 0.44 n/a 0.68 0.84 0.82
All 0.95 0.52 0.85 0.70 0.83 0.83
All 0.96 0.57 0.91 0.72 0.84 0.79
Impact-rev. 0.93 0.54 0.90 0.70 0.84 0.80
Impact-rev. 0.96 0.58 0.93 0.68 0.83 0.83
Table 13: Ablation results for Qwen and Gemini - Faithfulness: Faith., Coverage: Cov., Citation Year Compliance: Cyc., Specificity: Spec., Insightfulness: Insi., Trend Awareness: Trend. These results show that adding impact-revealing contexts and their intents has an improvement on almost all metrics.
Paper: Hierarchically Classifying Documents Using Very Few Words, ICML, 1997.
Qwen Gemini
Impact period: 1998-2004 Initial Adoption and Methodological Foundations: The paper laid the groundwork for hierarchical text classification by introducing a method that effectively classifies documents using very few words.. The initial adoption of this methodology was seen in several studies that built upon or compared their methods against the Koller and Sahami model. For example, the work by TaxE (2004) highlighted the challenge of comparing different hierarchical classification approaches due to the lack of a common dataset, underscoring the need for standardized evaluation metrics. Additionally, early works in 1998-2000 focused on extending and validating the use of hierarchical structures in text classification.
Impact period: 2005-2010 Extension to Various Domains and Techniques: During this period, the methodologies proposed in the paper were extended and applied to various domains, including medical vocabularies and web taxonomies. The work by Koller and Sahami was influential in demonstrating the effectiveness of hierarchical classification, which led to the development of new techniques such as the use of word embeddings and unsupervised learning. Research during this period also explored the integration of hierarchical structures with other machine learning algorithms like SVMs and neural networks. For instance, a 2007 study by Chakrabarti et al. drew parallels with Koller and Sahami’s approach to attribute selection, while another paper in 2011 critiqued existing feature selection methodologies, suggesting improvements for hierarchical classification.
Impact period: 2011-2020 Advanced Techniques and Real-World Applications: In this period, the paper’s influence continued to grow as researchers developed more advanced techniques for hierarchical text classification. The work by Koller and Sahami served as a foundational reference for numerous studies that sought to improve the robustness and efficiency of classification models. Specifically, research in 2018 explored the use of word embeddings for hierarchical text classification, building on the initial idea of leveraging hierarchical structures. Additionally, real-world applications of hierarchical classification were demonstrated in areas such as protein function prediction and image classification. The impact of the paper was also seen in the development of testbeds like TaxE, which aimed to standardize the evaluation of hierarchical classifiers.
Impact period: 1998-2005 Pioneering Hierarchical Text Classification and Feature Selection: The paper rapidly established itself as a seminal work in text classification. The paper was widely recognized for introducing and validating the use of class hierarchies to improve classification efficiency and accuracy over traditional flat models (e.g., [15, 16]). Researchers often cited it as the first proper study on hierarchical text categorization (e.g., [12]) and a primary reference for feature selection strategies in high-dimensional text data. However, early discussions also began to identify potential limitations, such as susceptibility to overfitting, the inability to recover from errors made high in the hierarchy, and the constraints of existing datasets for comprehensive evaluation (e.g., [12, 14]).
Impact period: 2006-2015 Methodological Evolution & Critical Analysis: This period solidified the paper’s status as a foundational reference for hierarchical classification, particularly its top-down approach, often referred to as the pachinko machine model. Researchers continued to build upon and compare against its principles, acknowledging its contributions to computational efficiency through techniques like subtree pruning (e.g., [33, 86]). However, there was a growing critical analysis of its inherent limitations, most notably the blocking issue or error propagation. The paper served as a baseline against which new algorithms were compared (e.g., [28, 92]), and its limitations, such as difficulties scaling to very large web taxonomies, were often highlighted to motivate new research (e.g., [8, 29]). The paper’s influence also extended to other domains like protein function prediction and image classification (e.g., [37, 56]).
Impact period: 2016-Present Enduring Relevance & Benchmark Status: The paper maintains its standing as a seminal and foundational benchmark for hierarchical text classification. Its principles are still acknowledged, even as the field evolves with advanced techniques like deep learning (e.g., [38, 41]). Crucially, the paper’s identified limitations, such as the propagation of errors in deep hierarchies, are frequently highlighted as motivations for current research to develop more robust and scalable solutions, particularly in areas like hierarchical multi-label classification and for handling large numbers of classes (e.g., [0, 10]).
Table 14: Examples of impact summaries generated by Qwen and Gemini - Part 1.
Paper: Comparability of telephone and face-to-face interviews in assessing axis I and II disorders,
The American Journal of Psychiatry, 1997.
Qwen Gemini
Impact period: 1997-2005 Validation of Telephone Assessments: The study provided empirical evidence that telephone interviews could yield results comparable to face-to-face interviews, thereby expanding the methods available for clinical and epidemiological research. This validation was particularly important for large-scale studies and surveys where face-to-face interactions were logistically challenging or cost-prohibitive. The study’s findings supported the reliability of tools such as the Personality Disorder Examination (PDE) and the Composite International Diagnostic Interview (CIDI) when administered via telephone, enhancing their utility in diverse settings.
Impact period: 2006-2012 Equivalence and Reliability in Mental Health Research: The influence of this paper extended into a broader validation of telephone assessments for various psychiatric disorders. Research during this period consistently cited the study’s findings to support the equivalence and reliability of telephone interviews compared to face-to-face methods. This period saw the application of these validated methods in numerous mental health studies, including those on major depressive disorder, anxiety disorders, and substance abuse disorders. The equivalence of telephone assessments was particularly valuable in longitudinal studies and clinical trials, where maintaining consistent and high-quality data collection was essential. Tools such as the SCID (Structured Clinical Interview for DSM-IV) were commonly used in these applications.
Impact period: 2013-2023 Practical Applications and Cost-Effectiveness: The practical applications of telephone interviews in mental health research and clinical practice became more prominent. The validated equivalence of telephone assessments led to increased adoption in various settings, including telehealth services, crisis intervention, and routine mental health care. This period also emphasized the cost-effectiveness of telephone assessments, particularly in large-scale surveys and during global events like the COVID-19 pandemic. The practicality and efficiency of telephone interviews were highlighted, making them a preferred method for data collection, especially when face-to-face interactions were not feasible.
Impact period: 1997-2007 Establishing Comparability and Validity of Telephone Interviews: The paper provided crucial evidence demonstrating that telephone interviews yield comparable results to traditional face-to-face assessments for various disorders, including anxiety and major depressive disorders. Researchers frequently cited the paper to support the validity of using telephone methods for large-scale data collection in clinical and epidemiological studies. The paper also began to inform discussions on the practical advantages of telephone interviews, such as cost-efficiency, broader geographical coverage, and potential for higher response rates, while also highlighting the importance of interviewer-subject rapport and identifying early nuances like potential underreporting of sensitive information (e.g., substance abuse) and the comparability for specific diagnostic tools like the CIDI or SCID.
Impact period: 2008-2017 Broadened Acceptance and Practical Utility: Over this period, the paper continued to be cited as authoritative evidence for the validity and reliability of telephone interviews, extending their application across a wider range of mental health conditions, including PTSD and eating disorders. Its influence shifted from merely establishing comparability to emphasizing the practical benefits of telephone methods, such as reduced participant burden, improved subject retention in longitudinal studies, and enhanced logistical efficiency for large-scale research. While broadly affirming the equivalence of telephone and face-to-face methods, citations during this time also reflected ongoing discussions about specific methodological considerations, such as the potential for differences in reporting stigmatizing behaviors or the role of nonverbal cues.
Impact period: 2018-2024 Enduring Relevance and Standard Practice: In recent years, the paper’s impact has solidified its position as a foundational reference for the routine use of telephone interviews in clinical assessment. It is cited as evidence for the established reliability and validity of telephone-based diagnostic methods, supporting their application as a standard practice for diagnosing various mental health conditions and even for monitoring adverse events. The citations highlight the practicality and effectiveness of telephone interviews, especially in contexts requiring reduced burden or remote data collection, such as during the COVID-19 pandemic.
Table 15: Examples of impact summaries generated by Qwen and Gemini - Part 2.
Appendix H Details on Model Usage
See Table 16.
Task/Component Model Name Details
intent classification and generation gpt-4o-mini temp=0, version=2024-07-18, Input: 128,000, Output: 16,384
impact summary generation gpt-4o temp=0, version=2024-11-20, Input: 128,000, Output: 16,384
impact summary evaluation gpt-4o temp=0, version=2024-11-20, Input: 128,000, Output: 16,384
textual pattern generation (for groundtruth data extension) gpt-4o temp=0, version=2024-11-20, Input: 128,000, Output: 16,384
practical applications (citation context analysis) gpt-4o-mini temp=0, version=2024-07-18, Input: 128,000, Output: 16,384
practical applications (author-level summary generation) gpt-4o temp=0, version=2024-11-20, Input: 128,000, Output: 16,384
Table 16: Details about the LLMs used in our experiments.
@@ I.1 Equal citation count, different citing reasons
We select 10 diverse research topics (e.g., LLMs for code generation, cultural sensitivity in clinical psychology), and analyze 3 papers within each. We ensure that the topic-specific papers have close citation counts. In this use case, we demonstrate how fine-grained citation intent analysis provides a deeper understanding of research impact, moving beyond raw citation counts. Results are in Tables 24 and 25. We observe that some papers’ main impact is inspiring or motivating new work, some are used for their method or data, others are cited to point out remaining challenges and propose improvements, etc. For example, for papers on open-information extraction with citation counts of $\sim$ 200,  Cui et al. (2018) is frequently cited to motivate new work or highlight existing limitations, while  Gashteovski et al. (2017) and  Han et al. (2019) are cited for their method. In commonsense knowledge mining, only one of the 3 papers is often cited for its dataset Hwang et al. (2021), and even though the other 2 papers are cited to highlight limitations, the limitations are different, namely challenges with handling constraints and reasoning for one Davison et al. (2019) , but data noise and limited coverage for the other Speer and Havasi (2013). These examples show that even when citation counts are similar, citation intents reveal the specific citing reasons, shaping the perception of a paper’s true impact and assisting in the impact summary generation task.
@@ I.2 Author-level impact summaries
Table 19: Human and LLM correlation and agreement on evaluating results (Section 3.3). p-value: $\leq$0.001, $\leq$0.05.
@@
Table 21: Examples of impact summaries from the informativeness evaluation, baseline = no-knowledge variant, our method = impact-revealing variants.